SOHO SCIENCE REPORT No. 7

April 1994

Editors:

Vicente Domingo and Bernhard Fleck

Space Science Department of ESA — ESTEC P.O. Box 299, NL–2200 AG Noordwijk The Netherlands Fax: +31 1719 84698 e-mail: vdomingo@soho.estec.esa.nl bfleck@soho.estec.esa.nl

Contents

1. Planned SOHO Science Working Team (SWT) activities 2. Announced SOHO science related conferences

3. SOHO related bibliography (III)

1. Planned SOHO Science Working Team (SWT) activities

Date  Event  Location  Contact

1994

May 15  SOI science team meeting  Los Angeles, CA  P. Scherrer (Stanford)
June 15-17 SWT–12  Portsmouth, England V. Domingo (ESTEC)
June 28-29 SOWG–11  GSFC, MD  V. Domingo (ESTEC)
July 6-8  LASCO/EIT Science Meeting Orsay, France  J.-P. Delaboudiniere (IAS)
Sep. 5-9  CELIAS Data Workshop ’94  Tramelan, Switzerland P. Bochsler (Univ. Bern)
Oct. 3-5  SUMER/CDS science meeting GSFC, MD  S. Jordan (GSFC)
Oct. 9-12  SOI science team meeting  Fallen Leaf Lake, CA P. Scherrer (Stanford)
Nov. 17/18 VIRGO science team meeting Brussels  C. Froehlich (Davos)

2. Announced SOHO science related conferences

Date  Event  Location  Contact

1994

May 16-20  CESRA Workshop on Coronal Potsdam,  A. Krüger
Magnetic Energy Releases  Germany  akrueger@aip.de
May 16-20  Helio- and Astero-Seismology Los Angeles, CA Roger Ulrich (UCLA)
from the Earth and Space  rulrich@solar.stanford.edu
May 23-27  SPD/AGU annual meeting  Baltimore, MD J. Gosling (@solar)
May 31 - June 2 3rd IACG Workshop  Tokyo (ISAS),  V. Domingo (ESTEC)
Solar events and their mani-  Japan

festations in interplanetary

space and geospace

June 5-10 8th International Symposium Sendai, Japan A. Morioka, Tohoku Univ.
on Solar-Terrestrial Physics  Fax: +81 222 262 6332
(STEP)  or: thkstp(40799)::morioka
June 13-17 2nd SOLTIP Symposium Nakaminato, Takashi Watanabe
Japan  41945::watanabe

watanabe@stelab.nagoya-u.ac.jp

Fax: +81 292 27 8532

SOHO Science Report No. 7 April 1994 2

2. Announced SOHO science related conferences (cont.)

Date  Event  Location  Contact

1994

July 10-22  COSPAR XXXth  Hamburg, Germany  Fax: +33 1 40 50 9827
Plenary Meeting  29531::cospar
July 24-29 SPIE’s Annual Meeting: San Diego, CA spie@mom.spie.org

Ultraviolet Technology V

Aug. 17-24  IAU XXIInd General Assembly:  The Hague,  V. Gaizauskas or J.O. Stenflo
” 19  Joint Discussion 6: Sun and  The Netherlands  M. Huber (ESTEC)
Heliosphere: Challenges for  mhuber@estsa2.estec.esa.nl
solar-terrestrial physics, magneto-  Fax: +31 1719 84699

and hydrodynamics

Sep. 19-21 5th Int. Conf. on Radiometry Berlin, Germany J. Fischer, PTB Berlin
NEWRAD ’94  fischer@merlin.berlin.ptb.de

Fax: +49 30 3481 490

Sep. 19-23  15th Sac Peak Summer Workshop Sacramento Peak Obs. Jeff Kuhn or Matt Penn
IR Tools for Solar Astrophysics New Mexico, USA jkuhn@sunspot.noao.edu
What’s Next?  mpenn@sunspot.noao.edu
Sep. 26-30 3rd SOHO Workshop Estes Park, CO M. Dryer (NOAA/SEL)
Solar Dynamic Phenomena &  mdryer@solar.stanford.edu

Solar Wind Consequences

Oct. 20-23 JOSO Annual Meeting: Trieste, Italy P. Brandt (KIS, Freiburg)
WG 6: Coordination of SOHO  pnb@kis.uni-freiburg.de

and ground-based observations

Dec. 5-9  AGU Fall Meeting  San Francisco, CA

1995

Feb. 7-10  Yosemite Conference on  Yosemite Nat. Park, CA J. Burch and H. Waite
Solar System Plasma Physics  Fax: +1 210 522 3493
April 2-6 4th SOHO Workshop Monterey, CA thoeksema@solar.stanford.edu

Helioseismology

May 29 - Jun. 2 AGU Spring Meeting  Baltimore, MD
June 25-30 Solar Wind 8 Dana Point Resort, CA D.J. McComas

Los Alamos Nat. Laboratory Los Alamos, NM 87545, USA

Dec. 4-8  AGU Fall Meeting  San Francisco, CA

3. SOHO related bibliography (III)

The references given below continue a list of SOHO related publications which was started in SOHO Science Report No. 5. The list is based on the IRS database of ESRIN/ESA as well as on references and pre-/reprints sent to us. If you are aware of publications which you think should appear here, please send us references, abstracts or pre-/reprints. A LaTeX file containing the complete list of SOHO related publications can be obtained on request from bfleck@soho.estec.esa.nl.

Bochsler, P.; Moebius, E.: Prospects for future solar-wind missions. In: Scientific Requirements for Future

Solar-Physics Space Missions. ESA SP-1157, 1993, p. 43. Bruls, J.H.M.J.: The formation of helioseismology lines. IV - The Ni I 676.8 nm intercombination line. Astronomy

and Astrophys. 269, 1993, p. 509. Engelhardt, W.; Hartwig, H.: Ultra-clean high-precision mechanisms for the SUMER EUV spectrometer on

SOHO. In: The Fifth European Space Mechanisms and Tribology Symposium, ESA SP-334, 1993. Engelhardt, W.; Hartwig, H.; Poser, K.: The development of the SOHO SUMER mechanisms. In: The Fifth

European Space Mechanisms and Tribology Symposium, ESA SP-334, 1993. Gabriel, A.H. and the GOLF Team: Global oscillations at low frequency from the SOHO mission (GOLF).

Advances in Space Research, vol. 11, no. 4, 1991, p. 103.

SOHO Science Report No. 7 April 1994 3

Garcia, R., Regulo, C., Roca Cortes, T.: Supergranulation contribution to the noise background velocity spectrum. Proc. Vth IRIS workshop and GOLF 93 annual meeting. eds. T. Roca Cortes and N.Garcia. 1993. p. 39-42.

Gomez, F.; Jimenez, A.; Herreros, J.M.; Roca Cortes, T.: El Proyecto VIRGO en SOHO. Tribuna de Astronomia.

1993. No.88, p. 47-53. Gonzales, A.; Herreros, J.M.; Roca Cortes, T.; Sequeiros, J.: El Observatorio Solar Espacial SOHO. Revista

Espanola de Fisica. 1993. Vol.7, No.1, p. 11-17. Habbal, S.R., Esser, R.: On the derivation of empirical limits on the helium abundance in coronal holes below

1.5 R. Astrophys. Journal Letters, 421, 1994, L59. Herreros, J.M.; Roca Cortes, T.; Sanchez, M.: El Proyecto GOLF en SOHO. Tribuna de Astronomia. 1993.

No.87, p. 51-58. Highman, C.O.; Woolaway, S.M.; Belmont, K.L.: Design of a mirror pointing and occulting mechanism for the

Ultraviolet Coronagraph Spectrometer (UVCS). In: Acquisition, tracking, and pointing VI. Proceedings of the Meeting of the Society of Photo-Optical Instrumentation Engineers. Orlando, Fl, 1992. p. 443.

Hollandt, J.; Huber, M.C.E., Kuehne, M.: Hollow Cathode Transfer Standards for the Radiometric Calibration

of VUV Telescopes of the Solar and Heliospheric Observatory (SOHO). Metrologia 30, 381-388. 1993. Huber, M.C.E.: Solar physics within ESA’s planning. In: Scientific Requirements for Future Solar-Physics Space

Missions. ESA SP-1157, 1993, p. 81. Huber, M.C.E.; Pedersen, A.; Taylor, B.G.; Wenzel, K.-P.; Perry, M.A.: Report on the activities of Space Science

Department 1990 - mid 1992. ESA SP-1161, 1993. Huomo, H.; Hannula, H.; Viherkanto, K.: Periscope mechanism for SWAN instrument. In: The Fifth European

Space Mechanisms and Tribology Symposium, ESA SP-334, 1993. Lancho, M.; Escobar, J.; Martinez, A.; Marczyk, J.: GOLF mechanism. In: The Fifth European Space Mecha-

nisms and Tribology Symposium, ESA SP-334, 1993. Lemaire, P.: Prospects for improved instrumentation on future solar space missions. In: Scientific Requirements

for Future Solar-Physics Space Missions. ESA SP-1157, 1993, p. 67. Mondello, G.; Basile, G.; Cherubini, G.; Tonetti, N.: Width modulator mechanism for spectrometer entrance

slits. In: The Fifth European Space Mechanisms and Tribology Symposium, ESA SP-334, 1993. Ogawa, H.S.; McMullin, D.; Judge, D.L.; Korde, R.: Normal incidence spectrophotometer using high density

transmission grating technology and highly efficiency silicon photodiodes for absolute solar EUV irradiance measurements. Optical Engineering. vol.32, no.12, 1993, p. 3121.

Regulo, C.; Palle, P.L.; Roca Cortes, T.: Velocity modelling of active regions crossing the Sun’s disk. Proc. Vth

IRIS workshop and GOLF 93 annual meeting. eds. T. Roca Cortes and N.Garcia. 1993. p. 69-72. Roca Cortes, T.; Sequeiros, J.: La Mision Cientifica Espacial SOHO. Politica Cientifica. 1993. No.38, p. 39-45. Timothy, J.G.; Bergamini, P.; Berger, T.E.; Bumala, R.W.; Liu, S.; Martinez, T.; Slater, D.C.: The Stanford

MAMA detector characterization facility. In: EUV, X-ray, and gamma-ray instrumentation for astronomy IV. Proceedings of the Meeting of the Society of Photo-Optical Instrumentation Engineers. San Diego, CA, 1993. 1993, p.59-67.

Timothy, J.G.; Bergamini, P.; Berger, T.E.; Bumala, R.W.; Slater, D.C.: Performance characteristics of the MAMA detectors for the SUMER instrument on the SOHO mission. In: EUV, X-ray, and gamma-ray instrumentation for astronomy IV. Proceedings of the Meeting of the Society of Photo-Optical Instrumentation Engineers. San Diego, CA, 1993. 1993, p.69-76.

Ulrich, R.K.; Henney, C.J.; Schimpf, S.; Fossat, E.; Gelly, B.; Loudagh, S., Schmider, F.-X.; Palle, P.; Regulo, C.; Roca Cortes, T.; Sanchez, L.: Modeling of integrated sunlight velocity measurements: the effect of surface darkening by magnetic fields. Astronomy and Astrophysics 280, Dec. 1993, 268-281.

Vial, J.-C.: The Solar Corona from SOHO. Adv. Space Res. Vol.14, No.4, p. 181-194, 1994.